# A Benchmark for Multi-speaker Anonymization

### v1

@@ Abstract
Privacy-preserving voice protection approaches primarily suppress privacy-related information derived from paralinguistic attributes while preserving the linguistic content. Existing solutions focus particularly on single-speaker scenarios. However, they lack practicality for real-world applications, i.e., multi-speaker scenarios. In this paper, we present an initial attempt to provide a multi-speaker anonymization benchmark by defining the task and evaluation protocol, proposing benchmarking solutions, and discussing the privacy leakage of overlapping conversations. Specifically, ideal multi-speaker anonymization should preserve the number of speakers and the turn-taking structure of the conversation, ensuring accurate context conveyance while maintaining privacy. To achieve that, a cascaded system uses spectral-clustering-based speaker diarization to aggregate the speech of each speaker and disentanglement-based speaker anonymization with a selection-based speaker anonymizer to conceal speaker privacy and preserve speech content. Additionally, we propose two conversation-level speaker vector anonymization methods to improve the utility further. Both methods aim to make the original and corresponding pseudo-speaker identities of each speaker unlinkable while preserving or even improving the distinguishability among pseudo-speakers in a conversation. The first method minimizes the differential similarity across speaker pairs in the original and anonymized conversations to maintain original speaker relationships in the anonymized version. The other method minimizes the aggregated similarity across anonymized speakers to achieve better differentiation between speakers. Experiments conducted on both non-overlap simulated and real-world datasets demonstrate the effectiveness of the multi-speaker anonymization system with the proposed speaker anonymizers. Additionally, we analyzed overlapping speech regarding privacy leakage and provide potential solutions [^1] Code and audio samples are available at https://github.com/xiaoxiaomiao323/MSA.
@@ I Introduction
One mainstream idea in the academic community to solve this problem is to implement a user-centric voice protection solution on the raw datasets before data sharing. In recent years, efforts to protect voice privacy have primarily concentrated on techniques like noise addition [8], voice transformation [9], voice conversion [10, 11, 12, 13], and speech synthesis [14]. However, researchers perform these studies in diverse settings, making them incomparable.
@@ II Related Work on
This section reviews SSA, which serves as the foundation for this study, including the goals outlined in VPC [15, 16, 17], as well as traditional and advanced SSA approaches.
@@ II-B1 Single-speaker anonymization approaches from VPC
Following VPC baseline models, SSA approaches can be categorized into digital signal processing- and disentanglement-based methods. Digital signal processing methods perceptively modify speech to conceal the original speaker’s identity. For instance, VPC provided one baseline where the speaker attributes are altered using the McAdams coefficient to shift the spectral envelope [15, 16, 17]. However, this approach may distort the content, even though the speaker’s privacy is being protected, making it less effective than disentanglement-based methods [15, 16, 17].
Two hypotheses underlie disentanglement-based approaches. First, speech can be explicitly decomposed into content, speaker identity, and prosodic representations, where speaker identity is a time-invariant representation across the utterance, while content and prosody are time-varying. This enables the modification of content and other paralinguistic attributes independently. Second, the representation of speaker identity carries most of the speaker’s private information; modifying the original speaker representation can hide the majority of speaker identity information. The modified speaker embeddings are finally combined with the original content and prosody representations to generate anonymized speech.
Because disentanglement-based SSAs show superior effectiveness in preserving speaker privacy and maintaining utility, the majority of speaker anonymization studies have adopted a similar disentanglement-based framework, with improvements made from several perspectives.
@@ II-B2 Improved speech disentanglement
Several works argue that the disentanglement of linguistic information extracted from the content encoder and prosody still contains original speaker information. To address this, modifications are made to the content and prosody. For instance, [23] removed speaker information from pitch and linguistic information by introducing differentially private noise. [24] modified pitch and speech duration to remove the speaker identity information residual in pitch contours and their dynamics. [25, 26] enhanced the disentanglement by removing speaker information from the acoustic model using vector quantization. [35] proposed a neural audio codec-based approach, known for generating high-quality synthetic speech when combined with language models, effectively bottlenecking speaker-related information to enhance privacy protection.
@@ II-B3 Improved speaker anonymizer
A widely-used selection-based speaker anonymizer [20, 21] replaced an original speaker vector with a mean vector (pseudo-speaker vector) of a set of randomly-selected speaker vectors from an external pool of English speakers. Previous research [15, 16] has demonstrated that anonymized voices generated by selection-based anonymizers have limited variability due to the average operation on speaker vectors. There have been recent attempts to preserve speaker distinctiveness. One of the top SSA systems [27] submitted to VPC 2022 utilized a generative adversarial network (GAN) to generate artificial speaker embeddings. This GAN was trained to turn random noise into artificial speaker embeddings that follow a similar distribution as the original speaker vectors. Another interesting work [28] is based on an orthogonal householder neural network (OHNN), which rotates the original speaker vectors into anonymized speaker vectors while ensuring they maintain the distribution over the original speaker vector space to preserve the naturalness of the original speech. The most recent work [29] models a matrix related to speaker identity and transforms it into an anonymized singular value transformation-assisted matrix to conceal the original speaker identity, generating more natural and distinctive anonymized speech.
@@ II-B4 Flexible attribute anonymization
SSA, as defined by VPC, focuses solely on removing speaker identity from the original speech while leaving the linguistic content and other paralinguistic attributes unchanged like age, gender, emotion, and dialect. However, these paralinguistic attributes, which also contain sensitive personal information, could potentially disclose a speaker’s privacy, including geographical background, social identity, and health status [30, 31]. Researchers have explored techniques aimed at protecting voice privacy by concealing various privacy-related characteristics, such as age, gender, and accent from speech signals.
@@ II-B5 Language-robust speaker anonymization
Another aspect of research aims at a SSA solution that can be applied to speech for unseen language. Self-supervised learning (SSL)-based speaker anonymization has been proposed [32, 28, 30], which utilizes an SSL-based content encoder to extract general context representations regardless of the input speech language. The entire system requires no text labels or other language-specific resources, enabling the system to anonymize speech data from unseen languages.
Applying SSA techniques directly in conversations would convert multi-speaker interactions into a single pseudo-speaker, thereby completely erasing the turn-taking information crucial for MSA. However, most SSAs can be extended to handle the multi-speaker conditions with additional components. In this work, we select SSL-based SSA as the backbone, as it has been verified to maintain good intelligibility, naturalness, and applicability across multiple languages [32, 33, 28]. This will be detailed in Section IV-B.
@@ IV A Cascaded Multi-speaker
The original conversation speech is fed into the SD module to generate rich transcription time-marked (RTTM) information for each speaker. The RTTM data is crucial as it serves as a foundation for aggregating the speech of each speaker and reconstructing the conversation sequentially. Subsequently, the individual single-speaker speech aggregated on the basis of RTTM data is anonymized to the same pseudo-speaker, while the speech from different speakers is anonymized into distinct pseudo-speakers222speaker-level anonymization. Notably, the background audio remains unaltered to preserve the authenticity and realism of the conversation. In this work, the widely-used spectral-based SD and the SSL-based SSA approach with selection-based anonymizer are chosen to establish a basic MSA framework [^3] Note that this pipeline is not limited to a specific SD and speaker anonymization method.. The details of each component are provided in the following.
@@ IV-A Speaker Diarization
The spectral clustering-based SD system involves multiple stages [39, 40], as illustrated in Fig. 2. First, a voice activity detection (VAD) system is used to filter out the non-speech regions. The active speech regions are then split into short fixed-length segments with a specific overlapping ratio. Subsequently, a pre-trained speaker embedding extractor is used to extract speaker vectors for each segment. Following this, a scoring backend, such as cosine scoring or probabilistic linear discriminant analysis (PLDA) [41, 42], is applied to compute similarity scores between pairs of segments. A clustering algorithm [43, 44, 45] is then used to assign a unique speaker label to each segment. Finally, the clustering results are summarized into the RTTM file, which contains the start time, duration, and speaker ID for each talking segment, providing the foundation for separating the speech of individual speakers.
@@ IV-B Speaker Anonymization
Speaker anonymizer: Hiding the original speaker’s identity for each speaker is crucial. Most works focus on modifying the original speaker embeddings using a speaker anonymizer, assuming that identity is mainly encoded in them. One commonly-used speaker anonymizer used in VPC baseline systems and based on an external pool is called selection-based anonymizer. It involves searching the $K$ farthest same-gender speaker vectors from an external speaker vector pool, then averaging $K^{*}<K$ randomly-selected ones as the anonymized speaker vector. Selecting the $K$ farthest speaker vectors as candidates to generate the anonymized speaker vector ensures that the anonymized speech does not sound like the original speaker, thus protecting the original speaker privacy. Averaging $K^{*}$ randomly-selected speaker vectors from the $K$ candidates helps prevent the leakage of speaker privacy for the pool speakers.
Anonymized speech generation: The anonymized speaker vector, original prosody, and content features are then fed into the neural vocoder HiFi-GAN [48] to generate individual anonymized segments. These segments are reconstructed into an anonymized conversation utilizing the temporal information provided in the RTTM file.
@@ 33       $\boldsymbol{x}_{\text{a}}^{i}=\boldsymbol{y}_{\text{a}}^{k}$
Algorithm 1 Anonymization function $f$
Given one conversation with $N$ speakers, let us denote their original speaker vectors as $\mathcal{X}_{\text{o}}=\{\boldsymbol{x}_{\text{o}}^{1},\ldots,\boldsymbol{x}_{\text{o}}^{N}\}$, where $\boldsymbol{x}_{\text{o}}^{n}\in\mathbb{R}^{D}$ is the $D$-dimensional segment-level speaker vector of the $n$-th speaker [^4] Each speaker has multiple segments in a conversation. These segments are aggregated into one single-speaker speech on the basis of diarization results, and then the speaker vector for each speaker is extracted, denoted $\boldsymbol{x}_{\text{o}}^{n}$.. Let the external pool with $P$ speaker vectors be $\mathcal{Y}_{\text{a}}=\{\boldsymbol{y}_{\text{a}}^{1},\ldots,\boldsymbol{y}_{\text{a}}^{P}\}$, where $\boldsymbol{y}_{\text{a}}^{p}\in\mathbb{R}^{D},\forall p\in[1,P]$ and $P\gg N$. The anonymization algorithm defines a function $f:\mathbb{R}^{D}\times\mathbb{R}^{D\times{P}}\rightarrow\mathbb{R}^{D}$, which produces an anonymized vector $\boldsymbol{x}_{\text{a}}^{n}=f(\boldsymbol{x}_{\text{o}}^{n},\mathcal{Y}_{a})$ for each $\boldsymbol{x}_{\text{o}}^{n}\in\mathcal{X}_{\text{o}}$. Note that the external pool $\mathcal{Y}_{a}$ is sourced from speakers different from the $N$ speakers to be anonymized [^5] To protect the identity of the speakers in the external pool, we average the ten most similar gender-consistent speaker vectors along with the original pool speaker vector itself to generate a replacement for this pool speaker vector. This is unlike the usual settings where the speaker vectors in the external pool are unanonymized..
Our goal is to design $f$ so that the anonymized speaker vectors well conceal the speakers’ identities and stay distinctive across speakers. Before explaining how the goodness of anonymization is measured, let us define a similarity matrix $\mathbf{S}^{x_{o}y_{a}}\in\mathbb{R}^{N\times{P}}$, where the element ${S}^{x_{o}y_{a}}_{i,j}$ is equal to the similarity between the original speaker vector $\boldsymbol{x}_{\text{o}}^{i}$ and the candidate anonymized vector $\boldsymbol{y}_{\text{a}}^{j}$. A common choice is to compute the cosine similarity, i.e.,
@@
After selecting the anonymized speaker vectors for each speaker, the similarities between the original and selected anonymized speaker vectors can be represented:
@@
$\mathcal{L}^{\text{DS}}(\mathcal{X}_{\text{o}},\mathcal{Y}_{\text{a}},f)$ $=\sum_{i=1}^{N}{S}^{x_{o}x_{a}}_{i,i}+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\left|{S}^{x_{a}x_{a}}_{i,j}-{S}^{x_{o}x_{o}}_{i,j}\right|,$ (6)
@@
$\mathcal{L}^{\text{AS}}(\mathcal{X}_{\text{o}},\mathcal{Y}_{\text{a}},f)$ $=\sum_{i=1}^{N}{S}^{x_{o}x_{a}}_{i,i}+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{S}^{x_{a}x_{a}}_{i,j}.$ (7)
To minimize the loss function (either Eq.(6) or Eq.(7)), we design an $f$ and describe it in the python-like Algorithm 1. Note that we perform gender-dependent anonymization in this implementation. This is done by separating the input $\mathcal{Y}_{\text{a}}$ and $\mathcal{X}_{\text{o}}$ into gender-dependent subsets and execute Algorithm 1 separately for female and male. This guarantees that the gender of each speaker remains the same before and after anonymization. When no gender annotation is available, a pre-trained gender recognition model predicts the gender.

### v2

@@ Abstract
Privacy-preserving voice protection approaches primarily suppress privacy-related information derived from paralinguistic attributes while preserving the linguistic content. Existing solutions focus particularly on single-speaker scenarios. However, they lack practicality for real-world applications, i.e., multi-speaker scenarios. In this paper, we present an initial attempt to provide a multi-speaker anonymization benchmark by defining the task and evaluation protocol, proposing benchmarking solutions, and discussing the privacy leakage of overlapping conversations. The proposed benchmark solutions are based on a cascaded system that integrates spectral-clustering-based speaker diarization and disentanglement-based speaker anonymization using a selection-based anonymizer. To improve utility, the benchmark solutions are further enhanced by two conversation-level speaker vector anonymization methods. The first method minimizes the differential similarity across speaker pairs in the original and anonymized conversations, which maintains original speaker relationships in the anonymized version. The other minimizes the aggregated similarity across anonymized speakers, which achieves better differentiation between speakers. Experiments conducted on both non-overlap simulated and real-world datasets demonstrate the effectiveness of the multi-speaker anonymization system with the proposed speaker anonymizers. Additionally, we analyzed overlapping speech regarding privacy leakage and provided potential solutions [^1] Code and audio samples are available at https://github.com/xiaoxiaomiao323/MSA, evaluation datasets can be download from https://zenodo.org/records/14249171..
@@ I Introduction
Regulations and the social awareness of the privacy issues have seen quite a few applications of the voice privacy protection techniques. For example, broadcasting companies have been using techniques to protect the identities of witness and whistle blowers in sensitive interviews [^2] https://www.nii.ac.jp/today/103/8.html (in Japanese). Similar techniques are used in medical sections to protect the privacy of the patients’ speech data [8]. Additionally, there are growing concerns in educational systems about protecting children’s speaker identities [^3] Privacy Guidelines for Children’s Online Safety while ensuring audio quality for analysis. Another application is to anonymize the speakers’ identities before publishing a speech database [9, 10]. This preserves the privacy of the original speakers while making the data available for downstream tasks.
One mainstream solution from the academic community is to implement a user-centric voice protection solution on the raw datasets before data sharing. In recent years, efforts to protect voice privacy have primarily concentrated on techniques like noise addition [11], voice transformation [12], voice conversion [13, 14, 15, 16], and speech synthesis [17]. However, researchers perform these studies in diverse settings, making them incomparable.
@@ II Related Work on
This section reviews SSA, which serves as the foundation for this study. It describes the goals outlined in VPC [18, 19, 20] as well as traditional and advanced SSA approaches.
@@ II-B1 Single-speaker anonymization approaches from VPC
SSA approaches can be categorized into digital signal processing (DSP)-based and disentanglement-based methods. A DSP-based method, for example the VPC baseline [37], conceals the perceived original speaker’s identity by warping the spectral envelope via McAdams coefficients. However, those methods distort the content more severely than disentanglement-based methods [18, 19, 20].
Two hypotheses underlie the disentanglement-based approaches. First, speech can be decomposed into content, speaker identity, and prosodic representations, where speaker identity is time-invariant across the utterance, while others are time-variant. Second, the speaker identity representation carries most of the speaker’s private information. By modifying the original speaker representation and using it and the original content and prosodic representations to reconstruct the anonymized speech, we expect to hide most of the speaker identity information in the original speech.
Because disentanglement-based SSAs show superior effectiveness in preserving speaker privacy and maintaining utility, the majority of speaker anonymization studies have adopted a similar disentanglement-based framework, with improvements made from several perspectives [^4] We summarize improvements into different subcategories, highlighting recently proposed novel and impactful methods. Some papers are cited multiple times for their contributions from various perspectives..
@@ II-B2 Improved speech disentanglement
Several works argue that disentangled content and prosodic representations still contain speaker information. To suppress leaked speaker information, various approaches have been proposed, for instance, adding differentially private noise to pitch and linguistic content [26], modifying pitch and speech duration [27], and applying vector quantization to content representation [28, 29]. Recently, a neural audio codec-based approach was proposed [38], effectively bottlenecking speaker-related information to enhance privacy protection. [39] employed a serial disentanglement approach to better separate global speaker identity representation, linguistic content, and prosody.
@@ II-B3 Improved speaker anonymizer
A widely-used selection-based speaker anonymizer [23, 24] replaced an original speaker vector with a mean vector (pseudo-speaker vector), which is an average of speaker vectors of randomly selected from an external pool of English speakers. Previous research [18, 19] has demonstrated that anonymized voices generated by selection-based anonymizers have limited variability due to the average operation. Facing the limitation, one of the top SSA submissions to VPC 2022 [30] utilized a generative adversarial network (GAN) to generate the anonymized speaker vectors from random noise. [31] uses an orthogonal householder neural network (OHNN) to rotate the original speaker vectors into anonymized ones, while ensuring the anonymized vectors follow the overall distribution of the original ones but do not overlap with them. Another work [32] applies singular value decomposition over a matrix composed from the speaker vectors. All these methods replace the selection-based anonymizer using linear or non-linear generative models.
@@ II-B4 Flexible attribute anonymization
SSA, as defined by VPC, focuses solely on removing speaker identity from the original speech. The linguistic content and other paralinguistic attributes remain unchanged, even though many of them, for example, age, gender, emotion, and dialect, could potentially disclose a speaker’s privacy, including geographical background, social identity, and health status [33, 34]. Researchers have explored techniques aimed at protecting voice privacy by concealing various privacy-related characteristics, such as age, gender, and accent from speech signals.
@@ II-B5 Language-robust speaker anonymization
Another area of research focuses on developing an SSA solution that can be applied to speech in unseen languages. Self-supervised learning (SSL)-based speaker anonymization has been proposed [35, 31, 33, 39], utilizing an SSL-based content encoder to extract general context representations regardless of the input speech’s language. The entire system requires no text labels or other language-specific resources, enabling it to anonymize speech data from unseen languages.
Despite advancements in SSA techniques, if we apply them directly to speech conversation data, they transform multi-speaker interactions into a single pseudo-speaker, erasing crucial turn-taking information essential for MSA. To preserve conversational dynamics while anonymizing individual speakers in MSA, one possible approach is to leverage SSA techniques as a foundation, adapting and enhancing them through additional modules to create a tailored solution. In this work, we select SSL-based SSA as the backbone, as it has been verified to maintain good intelligibility, naturalness, and applicability across multiple languages [35, 36, 31]. This will be detailed in Section IV-B.
@@ IV A Cascaded Multi-speaker
The original conversation speech is fed into the SD module to generate rich transcription time-marked (RTTM) information for each speaker. The RTTM data is crucial as it serves as a foundation for aggregating the speech of each speaker and reconstructing the conversation sequentially. Subsequently, the individual single-speaker speech aggregated on the basis of RTTM data is anonymized to the same pseudo-speaker, while the speech from different speakers is anonymized into distinct pseudo-speakers555speaker-level anonymization. Notably, the background audio remains unaltered to preserve the authenticity and realism of the conversation. In this work, the widely-used spectral-based SD and the SSL-based SSA approach with selection-based anonymizer are chosen to establish a basic MSA framework [^6] Note that this pipeline is not limited to a specific SD and speaker anonymization method.. The details of each component are provided in the following.
@@ IV-A Speaker Diarization
The spectral clustering-based SD system involves multiple stages [43, 44, 45], as illustrated in Fig. 2. First, a voice activity detection (VAD) system is used to filter out the non-speech regions. The active speech regions are then split into short fixed-length segments with a specific overlapping ratio. Subsequently, a pre-trained speaker embedding extractor is used to extract speaker vectors for each segment. Following this, a scoring backend, such as cosine scoring or probabilistic linear discriminant analysis (PLDA) [46, 47], is applied to compute similarity scores between pairs of segments. A clustering algorithm [48, 49, 50] is then used to assign a unique speaker label to each segment. Finally, the clustering results are summarized into the RTTM file, which contains the start time, duration, and speaker ID for each talking segment, providing the foundation for separating the speech of individual speakers.
Figure 3: Workflow of selection-based speaker anonymizer using an external speaker vector pool, adopted by VPC baseline systems. The input speaker vector $\boldsymbol{x}_{o}^{n}$ is anonymized by selecting $K$-farthest vectors in the pool $\mathcal{Y}_{\text{a}}=\{\boldsymbol{y}_{\text{a}}^{1},\ldots,\boldsymbol{y}_{\text{a}}^{P}\}$. The anonymized output $\boldsymbol{x}_{a}^{n}$ is set to be the average of the $K$-farthest vectors.
@@ IV-B Speaker Anonymization
Speaker anonymizer: Hiding the original speaker’s identity for each speaker is crucial. Most works focus on modifying the original speaker embeddings using a speaker anonymizer, assuming that identity is mainly encoded in them. Given one conversation with $N$ speakers, let us denote their original speaker vectors as $\mathcal{X}_{\text{o}}=\{\boldsymbol{x}_{\text{o}}^{1},\ldots,\boldsymbol{x}_{\text{o}}^{N}\}$, where $\boldsymbol{x}_{\text{o}}^{n}\in\mathbb{R}^{D}$ is the $D$-dimensional segment-level speaker vector of the $n$-th speaker [^7] Each speaker has multiple segments in a conversation. These segments are aggregated into one single-speaker speech on the basis of diarization results, and then the speaker vector for each speaker is extracted, denoted $\boldsymbol{x}_{\text{o}}^{n}$.. Let us then define an external pool of speaker vectors from $P$ speakers as $\mathcal{Y}_{\text{a}}=\{\boldsymbol{y}_{\text{a}}^{1},\ldots,\boldsymbol{y}_{\text{a}}^{P}\}$, where $\boldsymbol{y}_{\text{a}}^{p}\in\mathbb{R}^{D},\forall p\in[1,P]$. Each anonymized speaker vector $\boldsymbol{x}_{\text{a}}^{n}=f(\boldsymbol{x}_{\text{o}}^{n},\mathcal{Y}_{\text{a}})$ is obtained for speaker $n\in[1,N]$ by the speaker anonymizer $f$. One commonly-used speaker anonymizer in VPC relies on an external pool as shown in Fig. 3. The anonymization algorithm in VPC defines a function $f_{\text{VPC}}:\mathbb{R}^{D}\times\mathbb{R}^{D\times{P}}\rightarrow\mathbb{R}^{D}$, which produces an anonymized vector $\boldsymbol{x}_{\text{a}}^{n}=f(\boldsymbol{x}_{\text{o}}^{n},\mathcal{Y}_{a}),\forall{n}\in[1,N]$. Note that the external pool $\mathcal{Y}_{\text{a}}$ is sourced from speakers different from the $N$ speakers to be anonymized and $P\gg N$ [^8] To protect the identity of the speakers in the external pool, we average the ten most similar, gender-consistent speaker vectors along with the original pool speaker vector itself to generate a replacement for this pool speaker vector. This differs from the usual settings where the speaker vectors in the external pool are left unanonymized. The anonymizer function $f$ in the VPC processes each original speaker vector $\boldsymbol{x}_{\text{o}}^{n}$ individually by searching for the $K$ farthest, same-gender speaker vectors in the external speaker vector pool and averaging them to generate the anonymized speaker vector $\boldsymbol{x}_{\text{a}}^{n}$ sequentially. Anonymized speech generation: The anonymized speaker vector, original prosody, and content features are then fed into the neural vocoder HiFi-GAN [53] to generate individual anonymized segments. These segments are reconstructed into an anonymized conversation utilizing the temporal information provided in the RTTM file.
@@ 33       $\boldsymbol{x}_{\text{a}}^{i}=\boldsymbol{y}_{\text{a}}^{k}$
Algorithm 1 Anonymization function $f_{\text{MSA}}$
Different from SSA where the anonymizer $f_{\text{VPC}}$ process each $\boldsymbol{x}_{o}^{n}$ independently, we aim to design a more complicated $f_{\text{MSA}}:\mathbb{R}^{D\times{N}}\times\mathbb{R}^{D\times{P}}\rightarrow\mathbb{R}^{D\times{N}}$ so that the anonymized speaker vectors well conceal the speakers’ identities and stay distinctive across the $N$ speakers.
Before explaining how the goodness of anonymization is measured, let us define a similarity matrix $\mathbf{S}^{x_{o}y_{a}}\in\mathbb{R}^{N\times{P}}$, where the element ${S}^{x_{o}y_{a}}_{i,j}$ is equal to the similarity between the original speaker vector $\boldsymbol{x}_{\text{o}}^{i}$ and the candidate anonymized vector $\boldsymbol{y}_{\text{a}}^{j}$. A common choice is to compute the cosine similarity, i.e.,
@@
Suppose that the anonymized speaker vector $\boldsymbol{x}_{a}^{n}$ has been obtained for each speaker $n\in[1,N]$ by the multi-speaker anonymizer $\mathcal{X}_{a}=f_{\text{MSA}}(\mathcal{X}_{o},\mathcal{Y}_{a})$. The similarities between the original and selected anonymized speaker vectors can be represented:
@@
$\mathcal{L}^{\text{DS}}(\mathcal{X}_{\text{o}},\mathcal{Y}_{\text{a}},f_{\text{MSA}})$ $=\sum_{i=1}^{N}{S}^{x_{o}x_{a}}_{i,i}+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\left|{S}^{x_{a}x_{a}}_{i,j}-{S}^{x_{o}x_{o}}_{i,j}\right|,$ (6)
@@
$\mathcal{L}^{\text{AS}}(\mathcal{X}_{\text{o}},\mathcal{Y}_{\text{a}},f_{\text{MSA}})$ $=\sum_{i=1}^{N}{S}^{x_{o}x_{a}}_{i,i}+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{S}^{x_{a}x_{a}}_{i,j}.$ (7)
To minimize the loss function (either Eq.(6) or Eq.(7)), we design an $f_{\text{MSA}}$ and describe it in the python-like Algorithm 1. Note that we perform gender-dependent anonymization in this implementation. This is done by separating the input $\mathcal{Y}_{\text{a}}$ and $\mathcal{X}_{\text{o}}$ into gender-dependent subsets and execute Algorithm 1 separately for female and male. This guarantees that the gender of each speaker remains the same before and after anonymization. When no gender annotation is available, a pre-trained gender recognition model predicts the gender.